# Code Drift: Towards Idempotent Neural Audio Codecs

### v1

@@ I Introduction
Idempotence – the stability of a codec’s decoded output, and thus its encoded representation, under multiple rounds of encoding and decoding – is a desirable property for codecs. This is because codecs with low idempotence significantly degrade media under multiple encodings, and may therefore be poorly suited for real-world use. Such multiple-encoding scenarios can arise when digital media is repeatedly compressed for transmission, easier storage, and retargeting for different platforms. The increasing adoption of neural codec-based generative models for audio postprocessing and editing [6, 7, 8] may also result in repeated encoding and decoding of an audia final versionles before they reaches listeners. While some signal-processing codecs explicitly address idempotence [9], the idempotence of neural audio codecs has not been studied in depth. Rather, existing research has focused on perceptual transparency – the preservation of output quality under a single encoding.
@@ II Related Work
Studies of Neural Audio Codec Representations: Recent works have evaluated the utility of neural audio codec representations for both discriminative and generative downstream tasks [22, 23]. Others have focused on identifying specific properties within these representations and understanding their impact on downstream performance. Lemercier et al. [24] find that the codebooks learned by Encodec [3] exhibit conditional dependence and show that downstream generative modeling performance can be improved by reducing this dependence via an additional training objective. Similarly, we evaluate one specific codec property (idempotence) and examine its relationship to downstream generative modeling performance.
@@ III Measuring Idempotence
Metrics: To measure the stability of output audio under multiple rounds of encoding, we use the full-reference metrics Scale-Invariant Signal-to-Distortion ratio (SI-SDR) [30] and Perceptual Evaluation of Speech Quality (PESQ) [31]. PESQ gives a “perceptually-weighted” estimate of similarity between the original audio and audio that has been encoded and decoded (up to 8kHz), while SI-SDR operates on the available bandwidth. Given the original audio input as a reference, codecs with high idempotence should produce reconstructions with high PESQ and SI-SDR scores even after multiple encoding iterations. To analyze drift in the internal representations of neural codecs after multiple encoding iterations, we compute the match rate between tokens at each separate codebook level of the encoded representation after $n$ and $(n+1)$ encoding iterations. This value ranges from $100\%$ when all tokens match to $0\%$ when none do. While this is a coarse measure that does not account for codebook size or latent similarity, it allows us to observe the convergence of certain codecs towards “stable tokenizations” over multiple iterations. We also measure codebook use (the entropy of encoded token distributions as a percentage of the maximum achievable codebook entropy) to study whether codecs converge to small subsets of tokens after multiple encoding iterations. Higher entropy indicates broader codebook use.

### v2

@@ I Introduction
Idempotence – the stability of a codec’s decoded output, and thus its encoded representation, under multiple rounds of encoding and decoding – is a desirable property for codecs. This is because codecs with low idempotence significantly degrade media under multiple encodings, and may therefore be poorly suited for real-world use. Such multiple-encoding scenarios can arise when digital media is repeatedly compressed for transmission, easier storage, and retargeting for different platforms. The increasing adoption of neural codec-based generative models for audio postprocessing and editing [6, 7, 8] may also result in repeated encoding and decoding of an audio file before a final version reaches listeners. While some signal-processing codecs explicitly address idempotence or “tandem coding” [9, 10, 11, 12], the idempotence of neural audio codecs has not been studied in depth. Rather, existing research has focused on perceptual transparency – the preservation of output quality under a single encoding.
@@ II Related Work
Studies of Neural Audio Codec Representations: Recent works have evaluated the utility of neural audio codec representations for both discriminative and generative downstream tasks [25, 26, 27, 28]. Others have focused on identifying specific properties within these representations and understanding their impact on downstream performance. Lemercier et al. [29] find that the codebooks learned by Encodec [3] exhibit conditional dependence and show that downstream generative modeling performance can be improved by reducing this dependence via an additional training objective. Similarly, we evaluate one specific codec property (idempotence) and examine its relationship to downstream generative modeling performance.
@@ III Measuring Idempotence
Metrics: To measure the stability of output audio under multiple rounds of encoding, we use the full-reference metrics Scale-Invariant Signal-to-Distortion ratio (SI-SDR) [35] and Perceptual Evaluation of Speech Quality (PESQ) [36]. PESQ estimates “perceptually-weighted” similarity between the original audio and encoded/decoded audio (up to 8kHz), while SI-SDR operates on the available bandwidth. Given the original audio input as a reference, codecs with high idempotence should produce reconstructions with high PESQ and SI-SDR scores even after multiple encoding iterations. To analyze drift in the internal representations of neural codecs after multiple encoding iterations, we compute the match rate between tokens at each separate codebook level of the encoded representation after $n$ and $(n+1)$ encoding iterations. This value ranges from $100\%$ when all tokens match to $0\%$ when none do. While this is a coarse measure that does not account for codebook size or latent similarity, it allows us to observe the convergence of certain codecs towards “stable tokenizations” over multiple iterations. We also measure codebook use (the entropy of encoded token distributions as a percentage of the maximum achievable codebook entropy) to study whether codecs converge to small subsets of tokens after multiple encoding iterations. Higher entropy indicates broader codebook use.